Most hydrocarbons are highly combustible. That is, they can react spontaneously with oxygen (O2), forming the products carbon dioxide, (CO2), and water, (H2O). What mass of carbon dioxide will be produced by the complete combustion of 9.53 g of methyl octane (CH3C8H17)?
Answer #1

Write the balanced chemical equation for the combustion of methyl octane, CH3C8H17 as below.

CH3C8H17 + 14 O2 --------> 9 CO2 + 10 H2O

As per the stoichiometric equation,

1 mole CH3C8H17 = 9 moles CO2.

The atomic masses are

C: 12.011 u

H: 1.008 u

O: 15.999 u

The gram molar masses are

CH3C8H17 = (1*12.011 + 3*1.008 + 8*12.011 + 17*1.008) g/mol = 128.259 g/mol.

CO2 = (1*12.011 + 2*15.999) g/mol = 44.009 g/mol.

Mole(s) of CH3C8H17 corresponding to 9.53 g CH3C8H17 = (9.53 g)/(128.259 g/mol)

= 0.07430 mole.

Mole(s) of CO2 produced as per the stoichiometric equation =

(0.07430 mole CH3C8H17)*(9 moles CO2)/(1 mole CH3C8H17)

= 0.6687 mole.

Mass of CO2 produced = (0.6687 mole)*(44.009 g/mol) = 29.4288 g ≈ 29.4 g (ans, correct to 3 sig. figs).